# StreetSurfGS: Scalable Urban Street Surface Reconstruction with Planar-based Gaussian Splatting

### v1

@@ II-A Large Scene Reconstruction
Significant advancements have been made in image-based large scene reconstruction over recent decades. Some studies [16, 17, 18, 19, 20, 21, 22] employ a structure-from-motion (SfM) pipeline to estimate camera poses and generate a sparse point cloud. Subsequent research [23, 24, 25, 26, 27, 28, 29, 30] focuses on producing dense point clouds or triangle meshes using multi-view stereo (MVS) techniques [31]. These methods try to estimate the depth map by matching feature points across different views. For instance, Li et al. [29] introduced a marker-less shape and motion capture technique for multi-view video sequences, while Su et al. [30] developed an uncertainty-guided network along with an uncertainty-aware loss function to address unsupervised perception of uncertainty.
@@ II-C Gaussian Splatting
In the realm of geometry reconstruction, NeuSG [71] integrates the previous NeRF-based surface reconstruction method NeuS [49] into the 3DGS framework to transfer surface properties to the 3DGS. SuGaR [12] extracts explicit meshes from the 3DGS representation by regularizing Gaussians over surfaces. Building on SuGaR, GaMeS [72] uses these explicit meshes as input, parameterizing Gaussian components through vertices, enabling real-time modification of Gaussians by altering mesh components during inference. 2DGS [73] simplifies the 3D volume into 2D oriented planar Gaussian disks, while GOF [74] introduces a Gaussian opacity field, identifying a level set to enable precise geometry extraction. PGSR [14] employs a planer-based Gaussian representation and proposes an unbiased depth rendering method. However, these methods primarily cater to object-centric camera trajectories and are not well-suited for urban street scenes, which are typically long and narrow. Furthermore, their approaches often encounter memory limitations when handling the extensive data characteristic of urban environments. In our work, we make the first attempt to conduct scalable urban street scene surface reconstruction using Gaussian Splatting.
III Preliminaries on 3D-GS
3D Gaussian Splatting [5] explicitly models the scene using anisotropic 3D Gaussian primitives and generates images via rasterizing projected 2D Gaussians, achieving SOTA rendering performance in novel view synthesis task within neural scene representations. Each 3D Gaussian $G(x)$ is parameterized by a center position $\mu\in\mathbb{R}^{3}$ and the covariance $\Sigma\in\mathbb{R}^{3\times 3}$:
$G(x)=e^{-\frac{1}{2}(x-\mu)^{T}\Sigma^{-1}(x-\mu)},$ (1)
where $x$ is an arbitrary position within the scene, $\Sigma$ is parameterized by a scaling matrix $S\in\mathbb{R}^{3}$ and rotation matrix $R\in\mathbb{R}^{3\times 3}$ with $RSS^{T}R^{T}$. For rendering, opacity $\sigma\in\mathbb{R}$ and color feature $F\in\mathbb{R}^{C}$ are represented using spherical harmonics (SH) to model view-dependent color $c\in\mathbb{R}^{3}$. The 3D Gaussians $G(x)$ are projected onto the image plane as 2D Gaussians $G^{\prime}(x^{\prime})$ [75]. A tile-based rasterizer efficiently sorts them in front-to-back depth order and employs $\alpha$-blending:
$C\left(x^{\prime}\right)=\sum_{i\in N}T_{i}c_{i}\sigma_{i},\quad\sigma_{i}=\alpha_{i}G_{i}^{\prime}\left(x^{\prime}\right),$ (2)
where $x^{\prime}$ is the queried pixel, $N$ represents the number of sorted 2D Gaussians binded with that pixel, and $T$ denotes the transmittance as $\prod_{j=1}^{i-1}\left(1-\sigma_{j}\right)$.
@@ IV-A Planar-based Octree Gaussian Splatting Representation
Our approach begins with a sparse point cloud $\mathbf{P}$ generated using structure-from-motion (SfM) with COLMAP [21]. Following the methodologies described in [5, 9], we approximate an axis-aligned bounding box for the scene and partition the sparse point cloud via an octree-based scheme. Each input camera view is associated with its corresponding anchor Gaussians, optimizing the octree structure’s utility. The depth range bounding pairs $(i,j)$ are estimated as $(\hat{d}_{\text{min}},\hat{d}_{\text{max}})$ using the $0.05$ and $0.95$ quantiles from the depth distribution $\mathbf{D}=\{d_{i,j}\}$, incorporating scaling factors to accommodate varying camera intrinsics. The number of octree layers $K$ is determined by:
$K=\left\lfloor\log_{2}\left(\frac{\hat{d}_{\text{max}}}{\hat{d}_{\text{min}}}\right)\right\rfloor+1.$ (3)
The centers of the occupied voxels $\mathbf{V}\in\mathbb{R}^{M\times 3}$ are computed as:
$\mathbf{V}=\left\{\left\lfloor\frac{\mathbf{P}}{\epsilon}\right\rceil\cdot\epsilon,\ldots,\left\lfloor\frac{\mathbf{P}}{\epsilon/2^{K-1}}\right\rceil\cdot\epsilon/2^{K-1}\right\},$ (4)
where the base voxel size $\epsilon$ corresponds to the coarsest octree layer. The level of detail (LOD) for each anchor Gaussian is recorded in $\mathbf{L}\in\mathbb{R}^{M\times 1}$, forming a multi-level representation of the scene.
Upon structuring the octree and identifying the voxel centers $\mathbf{V}$, we convert the 3D Gaussians into 2D flat Gaussians to more accurately capture the surface geometry of the scene. Using the planar-based octree Gaussian splatting representation, we render depth and normal maps for subsequent regularizations, ensuring a precise and efficient urban scene reconstruction.
@@ IV-B1 Smooth Regularization
$\mathbf{N}_{d}(\mathbf{p})=\frac{(\mathbf{P}_{2}-\mathbf{P}_{1})\times(\mathbf{P}_{4}-\mathbf{P}_{3})}{\|(\mathbf{P}_{2}-\mathbf{P}_{1})\times(\mathbf{P}_{4}-\mathbf{P}_{3})\|},$ (5)
where $\mathbf{P}_{1}$ to $\mathbf{P}_{4}$ correspond to the spatial positions of the four neighbors of point $\mathbf{p}$ in the depth map. We then impose consistency between these computed normals and those directly derived from the Gaussian representation:
@@ IV-B2 Multi-view Consistency Regularization
Neighbour camera matching. For each pose, we select two cameras from the four nearest viewpoints. Depth maps are generated, and key points from the current viewpoint are transformed into the nearest cameras’ coordinate systems. These points are projected back into the current view to measure pixel-wise differences. Additionally, we compute the normalized cross-correlation (NCC) between patches in the reference and nearest images to enforce photometric consistency.
@@
Progressive Training. We avoid optimizing anchor Gaussians simultaneously across all LOD levels to prevent insufficient supervision at higher levels. Inspired by previous progressive training strategies [76, 55, 52], we implement a coarse-to-fine optimization approach, starting training at lower LOD levels and progressively activating higher levels, so that we can organizes anchor points across the LOD spectrum.
@@ V-A Experimental Setting
KITTI-360 is a large-scale dataset captured in urban environments with a driving distance of around $73.7$ km. We follow [3] to select $5$ sequences from them to evaluate our method. These sequences are relatively short, allowing us to compare StreetSurfGS with all the baseline methods, such as 3DGS [5] and PGSR [14]. Following [3], for the evaluation of novel view synthesis, we select every second image in each sequence as the test set and train our model on the remaining images. For the evaluation of geometry reconstruction, we use all the images for training.
Waymo Open is another urban environment dataset. We follow [82] and [3] to select the $6$ sequences mainly containing static objects for our experiments. Following [3], we select every $10$th image in the sequences as the test set and take the remaining ones as the training set. We employ identical settings for both novel view synthesis and geometry reconstruction tasks.

### v2

@@ II-A Large Scene Reconstruction
Significant advancements have been made in image-based large scene reconstruction over recent decades. To get camera positions and create a sparse 3D point cloud, multiple studies [16, 17, 18, 19, 20, 21, 22] leverage a structure-from-motion (SfM) pipeline. Many works [23, 24, 25, 26, 27, 28, 29, 30] also focus on producing dense point clouds or triangle meshes using multi-view stereo (MVS) techniques [31]. These methods try to estimate the depth map by matching feature points across different views. For instance, Li et al. [29] introduce a marker-less shape and motion capture technique for multi-view video sequences, while Su et al. [30] developed an uncertainty-guided network along with an uncertainty-aware loss function to address unsupervised perception of uncertainty.
@@ II-C Gaussian Splatting
In the realm of geometry reconstruction, NeuSG [71] integrates the previous NeRF-based surface reconstruction method NeuS [49] into the 3DGS framework to transfer surface properties to the 3DGS. SuGaR [12] extracts explicit meshes from the 3DGS representation by regularizing Gaussians over surfaces. Building on SuGaR, GaMeS [72] uses these explicit meshes as input, parameterizing Gaussian components through vertices, enabling real-time modification of Gaussians by altering mesh components during inference. 2DGS [73] simplifies the 3D volume into 2D oriented planar Gaussian disks, while GOF [74] introduces a Gaussian opacity field, identifying a level set to enable precise geometry extraction. To maintain a trade-off between rendering quality and geometry reconstruction, GSDF [75] combines 3DGS and SDF into a dual-branch architecture. PGSR [14] employs a planer-based Gaussian representation and proposes an unbiased depth rendering method. However, these methods primarily cater to object-centric camera trajectories and are not well-suited for urban street scenes, which are typically long and narrow. Furthermore, their approaches often encounter memory limitations when handling the extensive data characteristic of urban environments. In our work, we make the first attempt to conduct scalable urban street scene surface reconstruction using Gaussian Splatting.
III Preliminaries on 3D-GS and Octree Representation
The method of 3D Gaussian Splatting [5] represents scenes by using stretched 3D Gaussian blobs and creates images by flattening them onto a 2D plane. This approach has achieved top-level performance in tasks involving novel view generation and neural scene representations. Each Gaussian in 3D space, denoted by $\mathbf{H}(\mathbf{x})$, is defined by a central point $\bm{\mu}\in\mathbb{R}^{3}$ and a covariance matrix $\bm{\Sigma}\in\mathbb{R}^{3\times 3}$, and is described mathematically by the following expression:
$\mathbf{H}(\mathbf{x})=\exp\left(-\frac{1}{2}(\mathbf{x}-\bm{\mu})^{T}\bm{\Sigma}^{-1}(\mathbf{x}-\bm{\mu})\right),$ (1)
where $\mathbf{x}$ represents a point within the scene, and $\bm{\Sigma}$ is the covariance matrix, which can be factorized as $\bm{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T}$, with $\mathbf{R}$ being a rotation matrix and $\mathbf{S}$ a diagonal matrix indicating the scaling. An optimized rasterization pipeline sorts these 2D Gaussians by depth and applies a blending technique based on their transparency. The color at each pixel on the 2D plane is calculated through:
$\mathbf{C}=\sum_{i=1}^{n}{T_{i}\alpha_{i}\mathbf{c}_{i}},\quad T_{i}=\prod_{j=1}^{i-1}\left(1-\alpha_{j}\right)$ (2)
where $n$ represents the number of 2D Gaussians contributing to that pixel, $T$ refers to the transmission factor, $\mathbf{c}_{i}$ is the spherical harmonics representation of the $i$-th Gaussian.
Subsequent Scaffold-GS [9] uses anchor points for 3D scene gaussian representation. Builing on this, Octree-GS [11] aligns these anchor points with level-of-detail (LOD) requirements to form a multi-level representation of the scene, which enhances rendering quality and increases rendering speed.
@@ IV-A Planar-based Octree Gaussian Splatting Representation
Our approach begins with a sparse point cloud generated using structure-from-motion (SfM) with COLMAP [21]. Following the methodologies described in  [9] and [11], we approximate an axis-aligned bounding box for the scene and partition the sparse point cloud via an octree-based scheme. Details of anchor initialization can be found in [11].
After structuring the octree and identifying the voxel centers, we convert the 3D Gaussians into 2D flat Gaussians to more accurately capture the surface geometry of the scene. Using the planar-based octree Gaussian splatting representation, we render depth and normal maps for subsequent regularizations, ensuring a precise and efficient urban scene reconstruction.
@@ IV-B1 Smooth Regularization
$\mathbf{N}_{d}(\mathbf{p})=\frac{(\mathbf{P}_{2}(\mathbf{p})-\mathbf{P}_{1}(\mathbf{p}))\times(\mathbf{P}_{4}(\mathbf{p})-\mathbf{P}_{3}(\mathbf{p}))}{\|(\mathbf{P}_{2}(\mathbf{p})-\mathbf{P}_{1}(\mathbf{p}))\times(\mathbf{P}_{4}(\mathbf{p})-\mathbf{P}_{3}(\mathbf{p}))\|},$ (3)
where $\mathbf{P}_{1}(\mathbf{p})$ to $\mathbf{P}_{4}(\mathbf{p})$ correspond to the spatial positions of the four neighbors of point $\mathbf{p}$ in the depth map. We then impose consistency between these computed normals and those directly derived from the Gaussian representation:
@@ IV-B2 Multi-view Consistency Regularization
Neighbour camera matching. We further expand the neighbour camera matching strategy used in PSGR [14]. For each pose, we select two cameras from the four nearest viewpoints. Depth maps are generated, and key points from the current viewpoint are transformed into the nearest cameras’ coordinate systems. These points are projected back into the current view to measure pixel-wise differences. Additionally, we compute the normalized cross-correlation (NCC) between patches in the reference and nearest images to enforce photometric consistency.
@@
Progressive Training. We avoid optimizing anchor Gaussians simultaneously across all LOD levels to prevent insufficient supervision at higher levels. Specifically, we implement a coarse-to-fine optimization approach, starting training at lower LOD levels and progressively activating higher levels, so that we can organizes anchor points across the LOD spectrum.
@@ V-A Experimental Setting
The KITTI-360 dataset, recorded across urban areas over a distance of approximately $73.7$ km, provides a large-scale dataset for our evaluation. Following the approach in [3], we select $5$ specific sequences for testing our method. Due to the relatively short length of these sequences, we are able to directly compare the performance of StreetSurfGS against several baseline methods, including 3DGS [5] and PGSR [14]. For novel view synthesis evaluation, we use every second frame as test data, training the model on the remaining images, as recommended by [3]. When assessing geometry reconstruction, all available frames are used for model training.
The Waymo Open dataset also represents urban scenes. Following the selection criteria in [81] and [3], we use $6$ sequences that primarily capture static objects. Consistent with [3], we designate every $10$th frame in these sequences as test data, while the remaining frames are used for training. The same experimental setup is applied to both novel view synthesis and geometry reconstruction tasks.